OpenStudy (anonymous):

Write a unique polynomial in factored form with real coeff. that meets the given cond: zeros: 0(multiplicity of 2), 3(multiplicity of 1), 2 + 3i (multiplicity of 1); leading coefficient of 1

OpenStudy (anonymous):

Is the answer f(x)=1(x^2)(x-3)(x^2+9)

OpenStudy (anonymous):

seems right to me, I don't know what leading coefficient means though

OpenStudy (anonymous):

oh means you just put a number in the front of the whole equation

OpenStudy (anonymous):

hmm, I don't think that the complex one is correct. \[x^2 + 9 = 0 \Rightarrow x = \pm i \cdot 3\]

OpenStudy (anonymous):

ohh really..

OpenStudy (anonymous):

would it be x^2-4x+13

OpenStudy (anonymous):

@Stiwan

OpenStudy (anonymous):

I figured it out, instead of (x² + 9) it should be (x² -4x + 13). (x²-4x+13) has the complex roots 2+3i and 2-3i

OpenStudy (anonymous):

This is how i got there: \[(x-z) \cdot (x-\overline{z}) = x^2 - (z+\overline{z})x + z\overline{z}\]Now with\[z + \overline{z} = 4, z\overline{z} = 13\] the aforesaid follows

OpenStudy (anonymous):

hm, yeah that's correct. somehow I didn't see your responses, sry